### v1

@@ B.6 Implementation Details
We use $400$ inducing points for all datasets as we found this a reasonable trade off between computation time and accuracy. We initialise our inducing point locations at a subset of the data inputs. We use a mutlilayer perception for the latent variable encoder, as described in section B.2, with hyperparameters listed in table 2. The encoder weights are initialised from a truncated normal distribution centred on 0 with a standard deviation of (2hidden layer $\sqrt{(\frac{2}{\text{hidden layer size}})}$ [He et al., 2015].
@@
We use a minibatch size of 128 to calculate the loss. We take 100 Monte Carlo samples to integrate out the latent variables $\mathbf{w}_{i}=\mathcal{N}(\mu_{i},\Sigma_{i})$.
Subproblem convergence
There are many different ways of measuring convergence. We assume a subproblem to have converged when the difference between the mean of the lower bound for the last $t_{conv}$ steps and the mean of the lower bound for the last $0.5t_{conv}$ steps is smaller than the standard deviation of the losses for the last $t_{conv}$. We set $t_{conv}=2000$.
Acyclic constraint convergence
We stop the constrained optimisation and move onto the cool-down phase when the $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})<10^{-8}$.
@@
Each optimisation step consists of two steps. First, we optimise the variational parameters using natural gradients [Salimbeni et al., 2018], with a step size of 0.1. Second, we optimise the Gaussian process hyperparameters using Adam [Kingma and Ba, 2014]. For the warm up phase, we use a learning rate of $0.05$, for the constrained optimisation phase we use a learning rate of $0.01$ if $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})>0.1$ else we use $0.005$. In the cool-down phase, we use a learning rate of $0.01$.
@@
We did as many random restarts as we had computational resources for. This means we did 3 restarts for the three-variable dataset, 5 restarts for the 10-variable synthetic ER and SF datasets, $15$ restarts for Sachs and between 2 and 5 restarts for the Syntren graphs.
Appendix C DGP-CDE Details
In section 5.1 we introduce the discrete Gaussian process conditional density estimator (DGP-CDE), which is used to enumerate every possible causal structure for the 3 variable case. For this, a separate DGP-CDE is trained for each of the 25 possible causal graphs, with the Gaussian process prior for each variable defined in eq. 8. The DGP-CDE differs from the CGP-CDE in that it only takes $\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)}$ as inputs, whereas the CGP-CDE takes $\mathbf{X}_{\neg i}$ as inputs and then learns the adjacency matrix.
As the adjacency matrix is fixed for the DGP-CDE there is no acyclic penalty, so the loss function is just the lower bound $\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\Lambda},\boldsymbol{\phi}):=\sum_{i=1}^{D}\mathcal{F}_{i}(q_{i},\boldsymbol{\Lambda}_{i},\phi_{i})$. When all 25 separate DGP-CDE models have been trained, the most likely graph is selected by choosing the model with the highest marginal likelihood.
C.1 Implementation Details
@@
As the DGP-CDE is limited in the number of variables it can model by the need to enumerate over all graphs, there is no need to make the lower bound scalable using the methods discussed in section B.2 [Hensman et al., 2013]. Instead, we use the collapsed version of the lower bound which can be calculated in closed form [Titsias, 2009].
@@
We use a sum of a linear kernel defined in eq. 13 and a squared exponential kernel defined in eq. 14. This kernel means the kernel expectations needed to calculate the lower bound can be computed in closed form.
@@
As the enumeration over graphs is computationally expensive, we use 200 inducing points which we found to be sufficient. As we are not using stochastic variational inference [Hensman et al., 2013] for the DGP-CDE, we can no longer mini-batch so don’t use a multilayer perceptron for the latent variables. Instead, we initialise the latent variable mean as $\mu_{i}=0.1\mathbf{x}_{i}$ and standard deviation is randomly sampled $\Sigma_{i}\sim\text{Uniform}(0,0.1)$.
@@
The use of the collapsed bound for the DGP-CDE requires a different optimisation schedule than for the CGP-CDE. This consists of a two part optimisation scheme. Due to the loss function being highly non-convex and suffering from local optima, the first step of the optimisation scheme uses Adam [Kingma and Ba, 2014] with a learning rate of $0.05$ to get into a good region of the decision space. Once the loss function reaches the value a noise model would have, the optimisation scheme switches to the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm to perform gradient descent for the final part of the optimisation.
@@
We did 10 random restarts for each DGP-CDE, selecting the one with the best lower bound for our final result. We analysed the effect of the random restarts by permuting the order of the random restarts and plotting, in fig. 4, the mean and standard deviation of the SHD, SID and F1111 across permutations as the number of restarts increases. From the plot, it is clear that as the number of random restarts increases, the mean value of the metrics, on average, improves. This is because more random restarts allow a more thorough exploration of the loss function. It also shows that a higher value of $\mathcal{L}_{\text{ELBO}}$ does lead to better causal structure discovery.
Figure 4: Effect of adding random restarts for the DGP-CDE. We record the metrics (SHD, SID and F1) for the restart with the lowest loss out of a growing set of restarts, starting with one restart and adding one restart at a time. We randomly permute the order in which the random restarts are added $50$ times and plot the mean and standard deviation for each of the metrics. This shows that on average increasing the number of random restarts improves performance.
Appendix D Data Standardisation
For additive noise models, it has been shown that when data is sampled from a simulated DAG and not normalised, the marginal variance tends to increase along the causal order [Reisach et al., 2021]. When this is the case, it is possible to, at least partially, determine the causal order by ranking the marginal variance of variables. Reisach et al. [2021] demonstrate that commonly used synthetic datasets often have this property meaning a simple baseline method based on variance sorting and regression can perform as well as some recent continuous structure learning algorithms. This variance can be manipulated by rescaling (such as changing measurement units), which would then give a different causal order. As such, results that take advantage of the variance may be sensitive to measurement scales. Such effects may also persist in real-world datasets due to the measurement scale used.
To get rid of these effects, we standardise all datasets before applying any of the causal discovery methods discussed in this paper. This means our results for the baselines don’t match the results reported in the original papers for similar datasets [Montagna et al., 2023, Lachapelle et al., 2019, Rolland et al., 2022] (note that some of these works also define SHD of an anti-causal edge as 1 where we define it as 2).
Appendix E Data Generation Details
@@ 3 Variable Synthetic Data:
For each of the six distinct causal structures for three variables (up to permutations of the variables) we generated five datasets, each with 1000 samples. The data was generated by fitting a GP-CDE with a sum of a linear kernel and one other kernel randomly selected out of Matérn12, Matérn32, Matérn52, squared exponential and rational quadratic. The kernel hyperparameters are sampled from $\boldsymbol{\sigma}_{i}\sim\text{Uniform}(\mathbf{1},\mathbf{100})$ and $\boldsymbol{\theta}_{i}\sim\text{Gamma}(\mathbf{1.5},\mathbf{1})$ while the other parameters are sampled as $\boldsymbol{\phi}_{i}\sim\text{Uniform}(\mathbf{0.01},\mathbf{1})$ and $\mathbf{w}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$. The data is then sampled from this GP-CDE.
10 Variable Synthetic Data:
For each of the ER [Erdos et al., 1960] and SF [Barabási and Albert, 1999] sampling schemes, we generate five random 10 node graphs, with an expected number of edges of $15$ to ensure the graph is not too sparse or dense. Data is then generated for each node $X_{i}$ using
$X_{i}:=f^{NN}_{i}(\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)},\epsilon_{i}),$
where $\mathrm{PA}_{\mathcal{G}}(i)$ are the parents of $X_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,1)$ is sampled from $\epsilon_{i}\sim\mathcal{N}(0,1)$. $f^{NN}$ is a randomly initialised neural network with two layers, 128 units and ReLU activation functions. For each graph, we sample 1000 data points.
@@ 10 Variable Synthetic Data:
Syntren is a synthetic data generator that approximates real transcriptional regulatory networks [Van den Bulcke et al., 2006]. Networks are selected from previously described transcriptional regulatory networks. Relationships between the genes are based on the Michaelis-Menten and Hill enzyme kinetic equations. The kinetic equations contain some biological noise and some lognormal noise is added on top. We use the data generated by Lachapelle et al. [Lachapelle et al., 2019]. This contains 10 datasets of 500 samples, each with 20 nodes. The data was generated using the syntren data generator [Van den Bulcke et al., 2006], using the E.coli network with the default parameters except for the probability for complex 2-regulator interactions which was set to one [Lachapelle et al., 2019].
Sachs Dataset:
The Sachs dataset is a real-world dataset containing $11$ nodes and $853$ observations measuring the expression levels of proteins [Sachs et al., 2005]. While Sachs et al. provide a consensus graph based on domain knowledge [Sachs et al., 2005], it contains cycles and there is some debate around its validity [Mooij et al., 2016]. This has led to causal discovery methods using different ground truth graphs, for example, Yu et al. use a ground truth graph with $20$ edges while Lachapelle et al. and Montagna et al. use a ground truth graph with $17$ edges [Yu et al., 2019, Lachapelle et al., 2019, Montagna et al., 2023]. Given the uncertainty around the correct causal graph, we use the 17 edge graph used by Lachapelle et al. [Lachapelle et al., 2019] for the results presented in this paper as it is commonly used. However, we found CGP-CDE outperformed, or performed similarly in some metrics, to the other baselines for other widely used ground truth graphs.
Appendix F Details about Baselines
For SCORE [Rolland et al., 2022] and NoGAM [Montagna et al., 2023] we use the implementation in the dodiscover package https://github.com/py-why/dodiscover. For CAM [Bühlmann et al., 2014], NOTEARS [Zheng et al., 2018], GraN-DAG and GraN-DAG++ [Lachapelle et al., 2019] we use the implementations in the GraN-DAG github repository https://github.com/kurowasan/GraN-DAG.
@@ Appendix F Details about Baselines
SCORE and NoGAM have very similar hyperparameters and for both methods, we use the default values ridge regression suggested by Montagna et al. which they tuned to minimise the generalisation error on the estimated residuals [Montagna et al., 2023]. We also use the default values for both methods for all the hyperparameters for the Stein gradient and Hessian estimators and CAM pruning step except for the cutoff value for the CAM pruning step which we set to $0.01$ as is standard for CAM and follows Montagna et al. [Montagna et al., 2023]. NoGAM has the additional hyperparameter of number of cross validation models, for which we use the preset value of five.
@@
For CAM we use the default preliminary neighbourhood selection and DAG pruning methods and a cutoff value of $0.001$ for the edge pruning as suggested by the authors [Bühlmann et al., 2014].
@@
For NOTEARS we again use the default values, including a threshold of $0.3$ as is suggested as a reasonable choice by Zheng et al. [Zheng et al., 2018]. We also use the default values suggested by Lachapelle et al. [Lachapelle et al., 2019] for the augmented Lagrangian.
GraN-DAG & GraN-DAG++:
GraN-DAG and GraN-DAG++ have the same hyperparameters, for which we use the values selected by the authors based on small scale experiments on a single dataset generated from a 10 node ER graph [Lachapelle et al., 2019]. This includes setting each neural network to have 10 units and leaky-ReLU activation functions. We use the same number of hidden layers for each dataset as the authors, who use two hidden layers for ER and SF graphs and one hidden layer for the Sachs and Syntren datasets. We run both these methods with preliminary neighbourhood search and pruning, as is shown to be effective in reducing overfitting [Lachapelle et al., 2019]. We also follow the authors in using a batch size of 64. For both GraN-DAG and GraN-DAG++ we use the preset values for the augmented Lagrangian [Lachapelle et al., 2019].
Appendix G Computational Resources
The experiments in this paper were run on 4090 GPUs which have 24GB of memory.

### v2

@@ C.6 Implementation Details
We use $400$ inducing points for all datasets as we found this a reasonable trade off between computation time and accuracy. We initialise our inducing point locations at a subset of the data inputs. We use a multilayer perception for the latent variable encoder, as described in Section C.2, with hyperparameters listed in Table 1. The encoder weights are initialised from a truncated normal distribution centred on 0 with a standard deviation of (2hidden layer $\sqrt{(\frac{2}{\text{hidden layer size}})}$ (He et al., 2015).
@@
We use a minibatch size of 256 to calculate the loss. We take 50 Monte Carlo samples to integrate out the latent variables $\mathbf{w}_{i}=\mathcal{N}(\mu_{i},\Sigma_{i})$.
Acyclic constraint penalty
The acyclic constraint penalty has three hyperparameters: the coefficient to linearly increase the weighting of the penalty term by, $\rho$, the threshold value for the acyclicity, $\tau$, and the number of power iterations used to approximate the eigenvectors for the gradients of the acyclic constraint. We tuned these variables using five datasets generated from $10$ variable ER1.51.51.51.5 graphs. We chose $50$ power iterations, as it led to better performance than smaller numbers of iterations and after $50$ we saw minimal improvement. For the $10$ variable ER1.51.51.51.5 datasets, we found $\rho=50$ and $\tau=0.005$ leads to the best performance across SHD, SID and F1, while still converging in a reasonable amount of time. However, as the number of variables scaled, this value of $\rho$ took prohibitively long, except for the Syntren dataset. This is because the ELBO scales with the number of variables, $\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\Lambda}_{i},\phi_{i}):=\sum_{i=1}^{D}\mathcal{L}_{\text{ELBO},i}(q_{i},\boldsymbol{\Lambda}_{i},\phi_{i})$. Therefore, for the $3$, $20$ and $50$ variable datasets we scale $\rho$ with the number of variables, such that $\rho=5D$. This ensures the relative magnitude of the ELBO and the acyclic penalty remains roughly the same, meaning the time for the acyclic constraint phase to converge is not dependent on the number of variables. We did not find scaling $\rho$ affected the performance, but it did speed up the acyclic phase.
@@
Each optimisation step consists of two steps. First, we optimise the parameters of the variational distribution of inducing points $q(\mathbf{u}_{i})=\mathcal{N}(\mathbf{u}_{i}|\boldsymbol{m}_{u,i},\boldsymbol{S}_{u,i})$ using natural gradients, which have been shown to significantly improve training time for variational Gaussian processes (Salimbeni et al., 2018). We linearly increase the natural gradient step size from $0.0001$ to $0.1$ for the first five iterations, and then use a step size of $0.1$. Second, we optimise the Gaussian process hyperparameters using Adam with a learning rate of $0.05$ (Kingma & Ba, 2014).
@@
We did as many random restarts as we had computational resources for. This means we did 2 restarts for the three-variable dataset, 2 restarts for each of the 50 variable datasets and the Syntren dataset, and 1 restart for the 20 variable datasets.
Appendix D DGP-CDE Details
In Section 6.1 we introduce the discrete Gaussian process conditional density estimator (DGP-CDE), which is used to enumerate every possible causal structure for the 3 variable case. For this, a separate DGP-CDE is trained for each of the 25 possible causal graphs, with the Gaussian process prior for each variable defined in Equation 9. The DGP-CDE differs from the CGP-CDE in that it only takes $\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)}$ as inputs, whereas the CGP-CDE takes $\mathbf{X}_{\neg i}$ as inputs and then learns the adjacency matrix.
As the adjacency matrix is fixed for the DGP-CDE there is no acyclic penalty, so the loss function is just the lower bound $\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\Lambda},\boldsymbol{\phi}):=\sum_{i=1}^{D}\mathcal{L}_{\text{ELBO},i}(q_{i},\boldsymbol{\Lambda}_{i},\phi_{i})$. When all 25 separate DGP-CDE models have been trained, the most likely graph is selected by choosing the model with the highest marginal likelihood.
D.1 Implementation Details
@@
As the DGP-CDE is limited in the number of variables it can model by the need to enumerate over all graphs, there is no need to make the lower bound scalable using the methods discussed in Section C.2 (Hensman et al., 2013). Instead, we use the collapsed version of the lower bound, which can be calculated in closed form (Titsias, 2009).
@@
We use a sum of a linear kernel defined in Equation 19 and a squared exponential kernel defined in Equation 20. This kernel means the kernel expectations needed to calculate the lower bound can be computed in closed form.
@@
As the enumeration over graphs is computationally expensive, we use 200 inducing points, which we found to be sufficient. As we are not using stochastic variational inference (Hensman et al., 2013) for the DGP-CDE, we can no longer mini-batch, so don’t use a multilayer perceptron for the latent variables. Instead, we initialise the latent variable mean as $\mu_{i}=0.1\mathbf{x}_{i}$ and standard deviation is randomly sampled $\Sigma_{i}\sim\text{Uniform}(0,0.1)$.
@@
The use of the collapsed bound for the DGP-CDE requires a different optimisation schedule than for the CGP-CDE. This consists of a two part optimisation scheme. Due to the loss function being highly non-convex and suffering from local optima, the first step of the optimisation scheme uses Adam (Kingma & Ba, 2014) with a learning rate of $0.05$ to get into a good region of the decision space. Once the loss function reaches the value a noise model would have, the optimisation scheme switches to the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm to perform gradient descent for the final part of the optimisation.
@@
We did 10 random restarts for each DGP-CDE, selecting the one with the best lower bound for our final result. We analysed the effect of the random restarts by permuting the order of the random restarts and plotting, in Figure 3, the mean and standard deviation of the SHD, SID and F1111 across permutations as the number of restarts increases. From the plot, it is clear that as the number of random restarts increases, the mean value of the metrics, on average, improves. This is because more random restarts allow a more thorough exploration of the loss function. It also shows that a higher value of $\mathcal{L}_{\text{ELBO}}$ does lead to better causal structure discovery.
Figure 3: Effect of adding random restarts for the DGP-CDE. We record the metrics (SHD, SID and F1) for the restart with the lowest loss out of a growing set of restarts, starting with one restart and adding one restart at a time. We randomly permute the order in which the random restarts are added $50$ times and plot the mean and standard deviation for each of the metrics. This shows that, on average, increasing the number of random restarts improves performance.
Appendix E Data Standardisation
For additive noise models, it has been shown that when data is sampled from a simulated DAG and not normalised, the marginal variance tends to increase along the causal order (Reisach et al., 2021). When this is the case, it is possible to, at least partially, determine the causal order by ranking the marginal variance of variables. Reisach et al. (Reisach et al., 2021) demonstrate that commonly used synthetic datasets often have this property, meaning a simple baseline method based on variance sorting and regression can perform as well as some recent continuous structure learning algorithms. This variance can be manipulated by rescaling (such as changing measurement units), which would then give a different causal order. As such, results that take advantage of the variance may be sensitive to measurement scales. Such effects may also persist in real-world datasets due to the measurement scale used.
To get rid of these effects, we standardise all datasets before applying any of the causal discovery methods discussed in this paper. For the synthetic datasets, we also normalise each variable during the data generating process. This has been shown to create graphs that are not Var-sortable or R2-sortable, meaning the causal order cannot be determined using variances and correlation artefacts (Ormaniec et al., 2024). This means our results for the baselines don’t match the results reported in the original papers for similar datasets (Montagna et al., 2023; Lachapelle et al., 2019; Rolland et al., 2022) (note that some of these works also define SHD of an anti-causal edge as 1 where we define it as 2).
Appendix F Data Generation Details
@@ 3 Variable Synthetic Data:
For each of the six distinct causal structures for three variables (up to permutations of the variables), we generated five datasets, each with 1000 samples. The data was generated by sampling a GP-CDE with a sum of a linear kernel and one other kernel randomly selected out of Matérn12, Matérn32, Matérn52, squared exponential and rational quadratic. The kernel hyperparameters are sampled from $\boldsymbol{\sigma}_{i}\sim\text{Uniform}(\mathbf{1},\mathbf{100})$ and $\boldsymbol{\theta}_{i}\sim\text{Gamma}(\mathbf{1.5},\mathbf{1})$ while the other parameters are sampled as $\boldsymbol{\phi}_{i}\sim\text{Uniform}(\mathbf{0.01},\mathbf{1})$ and $\mathbf{w}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$. The data is then sampled from this GP-CDE.
20 & 50 Variable Synthetic Data:
For both 20 and 50 nodes, we create synthetic graphs using the ER (Erdos et al., 1960) and SF (Barabási & Albert, 1999) sampling schemes, with expected edges of one and four to simulate sparse and dense graphs. For each graph type, we generate five random graphs. Data is then generated for each node $X_{i}$ using
$X_{i}:=f^{NN}_{i}(\mathbf{X}_{\mathrm{PA}_{\mathcal{G}}(i)},\epsilon_{i}),$ (31)
where $\mathrm{PA}_{\mathcal{G}}(i)$ are the parents of $X_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,1)$ is sampled from $\epsilon_{i}\sim\mathcal{N}(0,1)$. $f^{NN}$ is a randomly initialised neural network with two layers, 128 units and ReLU activation functions. For each graph, we sample 1000 data points. This data generating scheme ensures that the final data is generated from a complicated distribution which does not directly correspond to any of the models.
@@
Syntren is a synthetic data generator that approximates real transcriptional regulatory networks (Van den Bulcke et al., 2006). Networks are selected from previously described transcriptional regulatory networks. Relationships between the genes are based on the Michaelis-Menten and Hill enzyme kinetic equations. The kinetic equations contain some biological noise and some lognormal noise is added on top. We use the data generated by Lachapelle et al. (2019). This contains 10 datasets of 500 samples, each with 20 nodes. The data was generated using the syntren data generator (Van den Bulcke et al., 2006), using the E. coli network with the default parameters except for the probability for complex 2-regulator interactions, which was set to one (Lachapelle et al., 2019).
Appendix G Details about Baselines
@@
For SCORE (Rolland et al., 2022) and NoGAM (Montagna et al., 2023) we use the implementation in the dodiscover package https://github.com/py-why/dodiscover. SCORE and NoGAM have very similar hyperparameters, and for both methods, we use the values of ridge regression suggested by Montagna et al., which they tuned to minimise the generalisation error on the estimated residuals (Montagna et al., 2023). We also use the values suggested by the authors for both methods for all the hyperparameters for the Stein gradient, Hessian estimators and CAM pruning step except for the cutoff value for the CAM pruning step which we set to $0.01$ as is standard for CAM and follows Montagna et al. (Montagna et al., 2023). NoGAM has the additional hyperparameter of number of cross validation models, for which we use the preset value of five.
@@
For CAM, we preliminary neighbourhood selection and DAG pruning method with a cutoff value of $0.001$ for the edge pruning as suggested by the authors (Bühlmann et al., 2014). We used the implementation in the https://github.com/kurowasan/GraN-DAG repository.
@@
For NOTEARS, we used the implementation provided in https://github.com/azizilab/sdcd, which uses binary search to find the adjacency matrix threshold that gives the largest possible DAG. This procedure has been found to give better performance than the default threshold of $0.3$ in the original NOTEARS paper (Nazaret et al., 2024; Lopez et al., 2022).
SDCD:
Nazaret et al. did not tune their hyperparameters for every dataset, instead using the same hyperparameter values, which were found to perform well empirically, for all their experiments (Nazaret et al., 2024, Appendix C.2). We use these same hyperparameter values and their implementation provided in https://github.com/azizilab/sdcd.
DiBS:
We use the non-linear model provided by the authors in https://github.com/larslorch/dibs/. For graph priors, we use the ER prior for the ER graphs and SF prior for the SF graphs, Syntren and the three variable baseline. This is because SF more closely matches the structure of the Syntren and the three variable datasets. For the other hyperparameters we use the values fine-tuned by the authors, including an observational noise of 0.1 and 3000 optimisation steps. The authors fine-tuned different values for the bandwidths of the kernel and slope of the linear schedule depending on the number of variables. We use these values, which are listed in (Lorch et al., 2021, Appendix E.3). DiBS typically returns a posterior over DAGs using Stein variational gradient descent. When a single particle is used, this becomes the MAP estimate of the graph. Therefore, as we wish to compare to our method, which returns a MAP estimate, we use one particle for DiBS, and perform five random restarts, selecting the one with the highest marginal likelihood value.
Random:
For the three variable dataset, we include uniform random sampling of DAGs. To compute this, for each ground truth graph, we enumerate all 25 possible graphs and evaluate their performance metrics. This exhaustive enumeration ensures that our results approximate those obtained in the limiting case where the number of sampled DAGs tends to infinity.
Appendix H Code Availability & Computational Resources
Code and data to replicate the experiments in this paper can be found at https://github.com/Anish144/ContinuousBMSStructureLearning. The experiments in this paper were run on A100 and RTX 4090 GPUs.
Appendix I Additional Experiments
Figure 4: 20 variable ER4 graphs, with data generated from (a) an additive noise model (Gaussian process) and (b) a non-additive noise model (neural network) as described by Equation 31. Black dot is the average metric. Lower is better for SHD and SID, and higher is better for F1. Despite the CGP-CDE being more expressive than the CGP, it is still able to perform comparably for the ANM dataset, demonstrating the added flexibility of the CGP-CDE does not hinder the performance on datasets generated from restricted models.
I.1 20 Variable ER4 Comparison with Additive Noise Model
Our model is more expressive than previous methods, many of which rely on the additive noise assumption. Here, we answer the following question: If the dataset is generated from an ANM model, how does our more expressive model perform?
To test this, we introduce the CGP, which is the same as the CGP-CDE model (Section 4), except that it is restricted to an additive noise model. This is done by simply removing the latent variable $\mathbf{w}$ in Equation 8. The likelihood then is a Gaussian, hence an ANM model. Removal of the latent variable also means that inference is more accurate in this model than the CGP-CDE, however at the cost of a loss of flexibility. All the settings of the CGP model are exactly the same as in Appendix C.
Figure 4(a) shows the performance on five datasets generated from 20 variable ER4 graphs with Gaussian process functional relations. Thus, the datasets were generated from an additive noise model. It is worth noting that the CGP outperforms other methods that make the ANM assumption, such as CAM, SCORE, NOGAM, DiBS, and SDCD. Despite being more expressive than the CGP model, the CGP-CDE model performs on par with the CGP model. This shows that with the added flexibility of our model, we do not lose performance on datasets generated from restricted models. This is due to the Occam’s razor effect of Bayesian model selection, which allows for expressing simpler models when the data requires it (Rasmussen & Ghahramani, 2000).
For comparison, Figure 4(b) shows results for five 20 variable ER4 non-additive noise datasets generated using a neural network as described in Equation 31. Here, the CGP-CDE outperforms the CGP due to its more flexible functional assumptions. Also, the CGP outperforms the other ANM methods, showing the benefit of the Bayesian model selection approach even when restrictive functional assumptions are made.
Figure 5: 20 variable ER4 graphs with data generated for 5 graphs as described by Equation 31. The x axis is the number of samples per graph, the circular markers are the mean and the vertical lines show the standard error across graphs. Lower is better for SHD and SID, higher is better for F1.
I.2 20 Variable with Different Numbers of Samples Results
To investigate the effect of the number of samples on the performance of the models, we performed experiments on five $20$ variable ER4444 graphs, with $100$, $500$ and $1000$ random samples. These results are shown in Figure 5. The CGP-CDE generally outperforms the other methods for metrics and numbers of samples, except for with $100$ samples, where CAM performs slightly better and NoGAM and SCORE perform comparatively, however, the CGP-CDE outperforms these methods on SID and F1.
I.3 20 Variable Results
To show the CGP-CDE performs well with a smaller number of variables on a range of graph types and densities, we performed experiments on $20$ variable graphs with the following graph types and densities: ER1 (Figure 6(a)), ER4 (Figure 4(b)), SF1 (Figure 6(b)), and SF4 (Figure 6(c)). These results show the CGP-CDE consistently performs well, although it is outperformed on some metrics for the ER1 and SF1 graphs, particularly by NOGAM, which seems to perform better on sparser graphs, perhaps because it is an ordering based method. However, the CGP-CDE is more consistent in performance than NOGAM across all our experiments with varying number of variables and graph types.
Figure 6: (a) Erdos-Renyi (ER) graphs with one expected edge per variable. (b) Scale-Free (SF) graphs with one expected edge per variable. (d) SF graphs with four expected edges per variable. For results for Erdos-Renyi (ER) graphs with four expected edges per variable, see Figure 4(b). Black dot is the average metric. Lower is better for SHD and SID, higher is better for F1.
I.4 Full 50 Variable Results
To demonstrate the performance of the CGP-CDE on a range of graph types and densities, we performed experiments on 50 variable graphs with the following graph types and densities: ER1, ER4, SF1, and SF4. The results of these experiments can be seen in Figure 7. These results show the CGP-CDE consistently performs well. For the ER graphs, the CGP-CDE outperforms all the baselines on SHD, SID and F1. For the SF graphs, where a few variables cause the rest of the variables, some of the baselines outperform the CGP-CDE on SID. We hypothesise this is because the structure of the SF graphs benefits methods that use topological search, such as NOGAM and SCORE. However, the CGP-CDE greatly outperforms these methods on SHD and F1 for both the SF graph densities.
Figure 7: (a) Erdos-Renyi (ER) graphs with one expected edge per variable. (b) ER graphs with four expected edges per variable. (c) Scale-Free (SF) graphs with one expected edge per variable. (d) SF graphs with four expected edges per variable. Black dot is the average metric. Lower is better for SHD and SID, higher is better for F1.